OpenStudy (anonymous):

A train travels 42 kilometers in 15 min. How far will the train travel in 3 hours at that speed. A. 168km B. 336km C. 504km D. 672km

OpenStudy (anonymous):

@shadow13

OpenStudy (black_ninja123):

What do you think the answer is?

OpenStudy (manish00333):

The answer is C) Method 1: There are 60 minutes in an hour, so 4 x 15 minutes 3 hours = (15 x 4) x 3 = 15 x 12 So there are 12 blocks of 15 minutes in 3 hours. You can then do: 12 x 42 = 504 km The train travels 504km in 3 hours. Method 2: 42 km in 15 15 x 4 = 1 hour and 42 x 4 = 168 so, it travels 168km in an hour 168(km) x 3(hours) = 504km in 3 hours

OpenStudy (black_ninja123):

60*3=180 180*42=7560 7560/15=504 @manish00333 is correct

OpenStudy (mathstudent55):

\(15 ~min = 0.25 ~h\) \(\dfrac{42~ km}{15~min} = \dfrac{42~km}{0.25~h} = 168 \dfrac{km}{h} \) \(168 \dfrac{km}{\cancel h} \times 3~\cancel h = 504~km\)

OpenStudy (theraggedydoctor):

Don't just give out answers. @manish00333

OpenStudy (anonymous):

Thanks @mathstudent55 @black_ninja123@manish00333